Consider the following for a firm. Its stock price (P0) is at $50, its payout ratio (POR) is 0.4, its EPS1 is $2.00, and investor required return is 10%.. What is the percentage of capital gains?


1
Expert's answer
2022-01-23T15:40:58-0500

Payout ratio (POR) = 0.4

Investor required return (IRR) = 10%

Then, the percentage of capital gains (CG) is:


"CG = (1-POR)\times IRR"


"CG = (1 - 0.4)\times 0.10=0.6 \approx6 \%"